28th June 2006 Hoyle committee room 13.00 - 14:10 Present: DWE, MJI, JRL, PSB, MR, EGS Apologies: WJS, RGM, STH Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Report on VISAS meeting 4. Data archives update 5. Optical processing update 6. WFCAM news 7. VDFS news 8. AOB Minutes ------- 1. Actions from last meeting MJI has made no further progress investigating the merging of classification statistics - ongoing <<<< WJS has not contacted the VPO to confirm that we will get no more ETC-related digitized data from them - dropped (see next action) MJI emailed WJS to remind him about contacting VPO. However, JPE has informed us that he is now dealing with the issue. Thus we have dropped the above action. PSB said that JPE has acquired some mirror reflectivity data (but has not passed this on to us). JRL has not implemented the fix to the problem with the UKIRT archive (missing header information problem) - ongoing <<<< EGS has not updated the INT WFC web pages with the latest flat-field and fringe frames - ongoing <<<< STH it is not known if the data from the 2MASS Touchstone fields has been acquired (STH not present) - ongoing <<<< DWE have updated the hardware page (HTML and Plone) MJI PSB had a short meeting to finalise updates to DRLD JRL MJI DWE are still finalizing content of Plone site - ongoing <<<< EGS STH MR 2. Comments on WFAU minutes The Rant of the Week was noted concerning the UDS processing. CASU extends its sympathies. MJI noted that the latest minutes stated that our requested network test would be carried out after DR1 has been released. 3. Report on VISAS meeting JRL reported on the VISAS meeting he attended in Edinburgh on 20 June. The main activity at the meeting was horse-trading between the original proposals, which was very successful. The original set of proposals shows a 3 to 4 times oversubscription in the available observing time. By merging some of these proposals, the amount of oversubscription will be reduced. From the 10 surviving proposals at the start of the meeting, there now remain 6 (Magellanic Clouds, Galactic Bulge, VIDEO, VIKING, VHS, UDS). What JRL found worrying was the way many of these projects were intending to take data. This will make things difficult for us. There is a need for us to give clear advice on the observing recipes. This comment is aimed at the proposals that don't want to dither. Many are concerned at the overheads that the ETC reported, so believe they will maximize their science by not dithering. PSB said that it was very important to dither due to the way that the bad pixels are distributed. Also, astronomers may not be aware that the acceptance level of the number of bad pixels per chip is 5%. JRL said that we have asked to produce a crib sheet of observing strategies. It would seem that there is a need to reexamine the overheads in the ETC, especially with respect to parallelization eg. reading data to disk while slewing. PSB reported a phone call where JPE said that he would write a specification for the overhead algorithm. MJI was sceptical and said that overheads were never overestimated. It was also noted that there was a distinct lack of manpower considerations for data reductions beyond that provided by the CASU pipeline in these proposals. ESO is likely to pick up on this. The merged proposals will now be rewritten and submitted to the selection panel. 4. Data archives update - AAT, ING, WFCAM, UKIRT ING - EGS said that the data for 1996 had now been copied (ingested?). From 1996 to 2006 is complete excepting for data from the first half of 2005 around the transition between DVD and Internet transfer of data from La Palma. MJI and EGS will arrange with RWA the filling in of this gap and stress test some more DVD readers. <<<< WFCAM - MR said that tapes have been received up to 15 June. All the data up to the 7th of June have been MEFed and are available through the raw data archive. He is currently dealing with the last set of tapes, 8th-15th of June. JRL said that he had a nephew coming for a couple of weeks for work experience and that MEFing files could be a suitable task. UKIRT - progress here is awaiting the missing header information problem fix. 5. Optical processing update EGS said that he had been reducing some INT data for Dan Smith and Martin Jarvis. 6. WFCAM news JRL says that it's linear. PSB asked who was going to tell 2MASS? MR reported on the progress of 06A reductions. The first half of May has missing chips in some of its files (as reported in previous minutes), but the pipeline now deals with these correctly. A small number of files have been found with the incorrect offsets in the header. MR was asked to email Paul Hirst (CC Andy Adamson) concerning this and any other problem he found. <<<< Since 2 June, all chips are present in the files, there are still some full of nulls for chip 2. This is about 1% of files. Some MSBs have been split and this has caused problems due to the way the MSB has been continued. These MSBs are effectively unreduceable. MR said that these MSBs were PI data. MR also said that he has been rereducing some 05B data for UDS. It was noted that in a recent paper UDS have claimed that our pipeline is not optimal. MR and EGS said that they will look into this. <<<< MJI said that the additional air conditioning has been installed successfully. This is a 7kW unit and is overspec. Two computers/RAID arrays have been rack mounted, thus freeing up PSB's desk. It is intended that two further computers (apm18 and apm23) will go into the store room. cass03 has been retired. There are no immediate plans to purchase any new hardware. 7. VDFS news PSB said that he had a great visit to RAL and that he has returned with 270 files - all with headers. Currently, RAL are in the middle of intensive spot projection test of VIRCAM. The data obtained included linearity sequences, noise/gain runs, darks and 1 persistence run. It is unlikely that we will get any more data. There was no schedule news. PSB seemed happy (we don't believe it either!). JRL reported that the ESO end-to-end test had been very successful. We supplied simulated data to test out the V0.3 recipe modules. They also carried out some archive ingestion tests using real RAL data. There were some problems reported, but only 1 (out of 6) was of a CASU origin. PSB said that it would be good to have hardware attached to VIRCAM that was comparable to what would be used at Paranal. This could be used to test the real overheads that could then be fed into the ETC. He wondered if this test could be carried out on a warm VIRCAM ie. after all the recent tests have been completed. 8. AOB None Continuing Actions ------------------ MJI finish investigating merged classification statistics JRL implement fix to problem with UKIRT archive (missing header information problem) EGS update INT WFC web pages with latest flat-field and fringe frames STH acquire data from the 2MASS Touchstone fields DWE finalize content of Plone site by end of month EGS STH MR New Actions ----------- MJI arrange with RWA to fill in the 2005 gap in the ING archive EGS MR email Paul Hirst (CC Andy Adamson) concerning incorrect offsets in file headers (and other problems) EGS investigate UDS claims regarding pipeline not being optimal MR